OpenStudy (anonymous):

to which graph does the point (-1,-4) belong

OpenStudy (anonymous):

can someone help me please

OpenStudy (goten77):

ugh... needs more info

OpenStudy (anonymous):

thats is the whole question do you want the multiple answers

OpenStudy (goten77):

to which graph does the point (-1,-4) belong theres infinite answers

OpenStudy (anonymous):

a. 3x + y = 12 b. 3x + y = -12 c. 3x - y = 12 d. 3x - y = -12

OpenStudy (anonymous):

these are the answer you can choose from

OpenStudy (goten77):

k well that limits it down then

OpenStudy (anonymous):

wrong ones wait one sec

OpenStudy (anonymous):

these are the real answers a. y < -x + 4 b. y < -x - 5 c. y < 2x - 3 d. y < 5x + 1

OpenStudy (anonymous):

the others are wrong

OpenStudy (anonymous):

think you could help me with some other after this

OpenStudy (goten77):

hmm i was wondering .... well anyways this is how u do it... well it has lots of ways to be done but ied do it like this id just directly plug in x/y and if its a true statement then thats the answer ..

OpenStudy (anonymous):

i cant view the whole answer

OpenStudy (goten77):

ctrl then scroll with ur mouse ... so u can zoom if thats the problem

OpenStudy (anonymous):

just type the answer

OpenStudy (goten77):

it is